OpenStudy (anonymous):

Part 1: Explain, in complete sentences, the pattern of exponents on the imaginary number i. Part 2: What is the usefulness of knowing this pattern when performing operations on complex numbers?

OpenStudy (inkyvoyd):

HINT: i^0=1 i^1=i i^2=-1 i^3=-i i^4=1 i^5=i i^6=...

OpenStudy (inkyvoyd):

If you know, then you can easily simplify i^(n) if n is a whole number, right?